Logiweb codex of hello world

Up Help

hello world,

hello world

[hello world →^{pyk} “hello world”]

The pyk compiler, version 0.grue.20050502 by Klaus Grue GRD-2005-06-01.UTC:19:29:19.961731 = MJD-53522.TAI:19:29:51.961731 = LGT-4624370991961731e-6